OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)^2

OpenStudy (anonymous):

help please

OpenStudy (cwrw238):

everything in the bravket is taken to the power 2 the rule is (x^a)^b= x^ab

OpenStudy (anonymous):

so its 6x^6 y^4

OpenStudy (anonymous):

(the answers 3x)^6

OpenStudy (anonymous):

(3x)^6

OpenStudy (cwrw238):

3 ^2 = 9 and (x^3)^2 = x^6 so the term in x comes to 9x^6 what about term in y?

OpenStudy (cwrw238):

sorry i got it wrong first time

OpenStudy (cwrw238):

do u follow that?

OpenStudy (anonymous):

9x^6 would be the answer then

OpenStudy (cwrw238):

- that the term in x but theres also y^4 what is (y^4)^2

OpenStudy (anonymous):

y^8

OpenStudy (cwrw238):

right so complete answer is 9x^6y^8

OpenStudy (anonymous):

k thx

OpenStudy (cwrw238):

yw